### v2

@@ Abstract
We study a multi-agent contract design problem with moral hazard. In our model, each agent exerts costly effort towards an individual task at which it may either succeed or fail, and the principal, who wishes to encourage effort, has an exclusive-use budget that it can use to reward the agents. We first show that any optimal contract must distribute the entire budget among the successful agents. Moreover, every such contract is optimal for some objective function. Our main contribution is then to introduce a novel class of contracts, which we call Luce contracts, and show that there is always a Luce contract that is optimal. A (generic) Luce contract assigns weights to the agents and distributes the entire budget among the successful agents in proportion to their weights. Lastly, we characterize effort profiles that can be implemented by Luce contracts, and note that Luce contracts offer a desirable alternative for implementation over commonly studied contracts, like piece-rate and bonus-pool contracts, on account of their reward variance-minimizing property.
@@ 1 Introduction
The literature on multi-agent contract design has identified conditions under which contracts like piece-rate contracts or bonus-pool contracts are optimal. These studies, however, implicitly assume that the principal is unconstrained in its ability to make payments, leading to some potentially undesirable features of the optimal contracts, such as significant variability in total payments and the possibility of excessively large payments. In practice, the principal’s capacity to pay may be limited due to liquidity constraints or organizational policies. In this paper, we model the contract design problem in the presence of such budgetary constraints and introduce a novel class of contracts called Luce contracts, demonstrating their optimality in such environments and also highlighting their relevance in more standard contract design environments.
@@
In our model, each agent exerts effort towards an independent task, in which it may succeed or fail, and the principal can commit to a contract which defines the payment to each agent as a function of the set of agents who are successful. We capture the principal’s budgetary restrictions by assuming it has access to an exclusive-use budget: the total payment to the agents can never exceed the budget, and any unused portion of the budget provides no additional value to the principal. This assumption is appropriate in situations where, for example, the budget is provided to the principal as a restricted-use fund, or where the principal can only reward the agents with probability shares for an indivisible prize. The principal’s problem is then to find a contract, which describes how the budget is distributed among the agents for each outcome, so as to incentivize agents to exert higher effort towards being successful.
@@
We first show that, even though there might be multiple contracts that are optimal, any such contract must be successful-get-everything (SGE): it rewards only the successful agents, and additionally, it distributes the entire budget among them. In particular, contracts such as piece-rate or bonus-pool contracts, which under certain outcomes would distribute only a fraction of the budget to successful agents, can never be optimal in this setting. Intuitively, the principal should be able to improve upon contracts that leave money on the table by using that leftover money to increase the rewards of some successful agent and thereby incentivize that agent to exert more effort. However, equilibrium forces imply that there isn’t necessarily an obvious way to distribute this left-over budget and induce higher effort from all agents. Nevertheless, we show that there is always some way to do this, and this implies that any optimal contract must be SGE. We further show that any contract which is SGE is optimal for some monotone objective of the principal. Stated differently, the maximal set of implementable equilibria are characterized by equilibria of SGE contracts.
@@
Our main contribution is then to introduce a subclass of SGE contracts, called Luce contracts, and to show that there is always a Luce contract that is optimal. A Luce contract is defined by assigning weights to and a (weak) priority order over the set of agents, and it distributes the entire budget among the highest-priority successful agents in proportion to their weights. We show that Luce contracts are minimally sufficient to implement the maximal set, and thus, the principal can always optimize over Luce contracts. This result provides for a significant reduction in the complexity of the design problem, narrowing the search space from a $\Theta(n2^{n})$-dimensional space of complex and difficult-to-interpret contracts to an $(n-1)$-dimensional space of simple and easily interpretable contracts. We apply our results to the special case of two agents with quadratic costs, where our analysis uncovers an interesting robustness in the structure of the optimal contract to the heterogeneity in the agents’ cost parameters.
@@
There is a vast literature on principal-agent problems under moral hazard. In the canonical model with a single agent (Holmström [20], Grossman and Hart [16], Mirrlees [33]), the principal offers a wage contract that defines the agent’s payment as a function of its observed output, and then the agent chooses some unobserved action (effort) that determines the distribution over outputs. A key finding is that optimal contracts reward the agent for output realizations that are informative about the target level of effort (informativeness principle), and may therefore be non-monotone in output. There has since been significant work studying variants of this single-agent model incorporating flexible actions (Georgiadis, Ravid, and Szentes [14]), multiple tasks (Holmstrom and Milgrom [23], Bond and Gomes [5]), bounded payments (Jewitt, Kadan, and Swinkels [26]), combinatorial actions (Dütting, Ezra, Feldman, and Kesselheim [9]), and informationally robust design (Carroll [6], Zhang [38]). See Georgiadis [13], Holmström [22] and Lazear [28] for recent surveys of this literature.
Our paper contributes to the literature studying a principal contracting with multiple agents. In these settings, agents may be rewarded not only based on their absolute performance but also on their relative performance or even the joint performance of the group. When agents’ observed outputs are subject to a common shock or involve subjective evaluations, the optimal contract may involve pay for relative performance (Green and Stokey [15], Lazear and Rosen [29], Malcomson [32, 31], Mookherjee [34], Nalebuff and Stiglitz [35], Imhof and Kräkel [24]). And joint performance evaluation may be optimal in the presence of negatively correlated outputs (Fleckinger [12]), complementarities in production (Alchian and Demsetz [1]), incentives to help (Itoh [25]), and unknown production environment (Kambhampati [27]). Other related work on multi-agent contract design explores moral hazard in teams (Holmstrom [21], Winter [37], Battaglini [4], Babaioff, Feldman, Nisan, and Winter [2], Halac, Lipnowski, and Rappoport [19], Dai and Toikka [8]), computational challenges in finding optimal contracts (Castiglioni, Marchesi, and Gatti [7], Dütting, Ezra, Feldman, and Kesselheim [10], Ezra, Feldman, and Schlesinger [11]), and design problems tailored to more specific environments (Halac, Kartik, and Liu [18], Haggiag, Oren, and Segev [17], Baiman and Rajan [3]). In contrast to this literature, our work focuses on a principal contracting with multiple independent agents, where standard single-agent contracts would be optimal but are rendered infeasible due to budgetary constraints on the principal.
@@ 2 Model
There is a principal and $n$ risk-neutral agents. Each agent $i\in[n]$ participates in an independent task, in which it may succeed or fail. The agents choose how much effort to exert towards succeeding in their respective tasks, captured by their choice of success probabilities: if agent $i$ chooses success probability $p_{i}\in[0,1]$, it incurs a cost $c_{i}(p_{i})$ in doing so, where $c_{i}:[0,1]\to\mathbb{R}$ is differentiable, strictly increasing, and strictly convex with $c_{i}^{\prime}(0)=0$. The cost functions are assumed to be common knowledge.
The principal would like to incentivize the agents to maximize the probability that they succeed in their tasks. Formally, we assume that the principal’s preference over success probability profiles is represented by a continuous, strictly increasing objective function $V(p_{1},\dots,p_{n})$. For example, if the principal obtains a profit of $w_{i}$ if agent $i$ succeeds in its task, and it cares about maximizing expected profit, its preference would be represented by the objective function
@@
In order to incentivize the agents, the principal can design a contract mapping observed outcomes to a reward for each agent. We assume that the principal has an exclusive-use budget $B$ that can be used for the purpose of rewarding the agents, and the principal gets no value from unused portions of the budget. Formally, the principal commits to a contract which specifies the fraction of the budget each agent receives depending on the observed outcome.
@@
$u_{i}(p)=\mathbb{E}[f_{i}(S)\cdot B]-c_{i}(p_{i})=\sum_{S\subset[n]}{\Pr}^{[n]}_{p}(S)\cdot f_{i}(S)\cdot B-c_{i}(p_{i})$
@@
Observe that since agent $i$’s expected reward is linear in $p_{i}$ and its cost function is strictly convex, $u_{i}$ is strictly concave in $p_{i}$. It follows then that for any contract $f\in\mathcal{F}$, a pure-strategy Nash equilibrium exists.
We will denote by $E:\mathcal{F}\rightrightarrows[0,1]^{n}$ the equilibrium correspondence, so $E(f)$ is the set of equilibria for the contract $f$ and $E^{-1}(p)$ is the set of contracts for which $p$ is an equilibrium. We further denote by $\mathcal{E}$ the set of implementable profiles:
@@
The principal’s problem is to choose a contract $f$ that induces an equilibrium $p\in E(f)$ that maximizes its objective $V$. As is standard, we assume that when a contract has multiple equilibria, the principal can select its most preferred equilibrium. [^1] For an illustration of multiplicity, suppose $B\geq\sum_{i\in[n]}c_{i}^{\prime}(1)$ and consider the bonus-pool contract under which agent $i$’s reward is $c_{i}^{\prime}(1)$ if all agents succeed and 00 otherwise. For this contract, it is easy to verify that both $p=(1,1,\dots,1)$ and $p=(0,0,\dots,0)$ are Nash equilibria. That said, if the budget is sufficiently small and the cost functions are sufficiently convex, we can use the diagonally strict concavity condition of Rosen [36] to show that every contract has a unique equilibrium. Hence, the principal’s problem is to maximize its objective $V$ over the set of implementable profiles $\mathcal{E}$. As we show, $\mathcal{E}$ is compact, and since $V$ is strictly increasing, any maximizer of $V$ must be a maximal element of $\mathcal{E}$ with respect to the standard coordinatewise ordering. We denote by $\mathcal{P}$ the set of maximal equilibrium profiles in $\mathcal{E}$. Formally,
@@
Notice that if $B\geq\sum_{i\in[n]}c_{i}^{\prime}(1)$, the principal can induce any $q\in[0,1]^{n}$ using the following piece-rate contract:
@@
Indeed, agent $i$’s expected utility in this case is $p_{i}\cdot c_{i}^{\prime}(q_{i})-c_{i}(p_{i})$ which is uniquely maximized at $p_{i}=q_{i}$. It follows that if the budget is sufficiently large, then $\mathcal{E}=[0,1]^{n}$, $\mathcal{P}=\{(1,1,\dots,1)\}$, and the principal’s problem is trivial.
@@
Assumption 1 ensures that in equilibrium, no agent chooses $p_{i}=1$. As we discuss later, our results extend to situations with a larger budget as long as we restrict attention to equilibria that are interior.
For convenience, we further normalize the budget to $B=1$. Note that this is without loss of generality, since for any contract $f\in\mathcal{F}$, the game with budget $B$ and costs $c_{i}$ is strategically equivalent to the game with budget $1$ and costs $\frac{1}{B}\cdot c_{i}$.
Best response
For any contract $f\in\mathcal{F}$, $u_{i}(p)$ is strictly concave in $p_{i}$. It follows that for any profile $p_{-i}$, agent $i$ has a unique best response, which we denote by $b_{i}(f,p_{-i})$. To find this best response, notice that the marginal utility of agent $i$ at profile $p$ is
$\dfrac{\partial u_{i}(p)}{\partial p_{i}}=r_{i}(f,p_{-i})-c_{i}^{\prime}(p_{i}),$
where
$r_{i}(f,p_{-i})=\mathbb{E}[f_{i}(S)\,|\,i\in S]-\mathbb{E}[f_{i}(S)\,|\,i\notin S]=\sum_{S\subset[n]_{-i}}(f_{i}(S\cup\{i\})-f_{i}(S)){\Pr}^{[n]_{-i}}_{p_{-i}}(S)$
denotes agent $i$’s expected gain in reward from succeeding as compared to failing in her task. Since $u_{i}$ is strictly concave in $p_{i}$, the best response is given by the solution to the first order condition. Moreover, since $r_{i}(f,p_{-i})\leq\mathbb{E}[f_{i}(S)\,|\,i\in S]\leq 1$ and by Assumption 1, $c_{i}^{\prime}(1)>B=1$, this best response is strictly smaller than $1$. In particular, we get that $b_{i}(f,p_{-i})$ is the solution to the equation
$c_{i}^{\prime}(p_{i})=\max\{0,r_{i}(f,p_{-i})\}.$ (1)
It follows that for any contract $f\in\mathcal{F}$, $p$ is an equilibrium under $f$ if and only if $p_{i}=b_{i}(f,p_{-i})$ for all $i\in[n]$.
In this paper, we identify contracts whose equilibria characterize $\mathcal{P}$, and further provide a subclass of contracts that is actually sufficient to implement $\mathcal{P}$. We use these results to solve for some natural objectives of the principal, and also briefly discuss the intermediate budget case towards the end of the paper.
3 Characterization of optimal contracts
In this section, we identify a class of contracts whose equilibria characterize the set of maximal equilibria.
3.1 FGN contracts
A natural property that one might expect any reasonable contract must have is that agents who fail get no reward.

### v3

@@ Abstract
We study a multi-agent contract design problem with moral hazard. In our model, each agent exerts costly effort towards an individual task at which it may either succeed or fail, and the principal, who wishes to encourage effort, has an exclusive-use budget that it can use to reward the agents. A motivating application is crowdsourcing for innovation, where a fixed budget is provided to a crowdsourcing platform to use for rewarding participants based on their submissions. Our main contribution is to introduce a novel class of contracts, which we call Luce contracts, and show that there is always a Luce contract that is optimal. A (generic) Luce contract assigns weights to the agents and distributes the entire budget among the successful agents in proportion to their weights. Furthermore, we characterize effort profiles that can be implemented by Luce contracts and show that Luce contracts offer a way to mitigate the uncertainty in total payments compared to alternative contracts—such as piece-rate or bonus-pool contracts—suggesting their desirability even in environments without budget constraints.
@@ 1 Introduction
We study a contract design problem for a principal interacting with many independent agents, as is the case for example in crowdsourcing environments. While the principal could, in theory, treat each agent independently and offer an individually optimal piece-rate contract, this incentive scheme has several undesirable features, including significant variability in total payments and the risk of excessively large payouts. In practice, contracts in such environments are often not piece-rate but instead structured as competitions, where the principal sets aside a fixed budget and commits to distributing the budget among a subset of well-performing agents. For example, crowdsourcing platforms like Wazoku and HeroX invite organizations to pose their unsolved problems as Challenges to a crowd of solvers, outlining criterion for a submission to be successful and specifying a budget for awarding successful submissions. In this paper, we model the contract design problem in the presence of budgetary constraints, and further, formalize the sense in which the competition structure of these contracts helps mitigate the uncertainty in total payments.
@@
In our model, each agent exerts costly effort towards a task, in which it may independently succeed or fail, and the principal, who is budget-constrained, can commit to a contract which defines the payment to each agent as a function of the set of agents who are successful. For convenience, we assume that the principal’s budget is exclusive-use, meaning that any unused portion of the budget provides no additional value to the principal. The assumption is appropriate when the budget is provided as a restricted-use fund or when agents are rewarded with probability shares for an indivisible prize. It is also reasonable in environments where the budget is small in comparison to the benefit associated with successful submissions, as might be the case in crowdsourcing for innovation. The principal’s problem is then to find a contract, which describes how the budget is distributed among the agents for each outcome, so as to incentivize agents to exert higher effort towards being successful.
@@
We first show that, even though there might be multiple optimal contracts, any such contract must be successful-get-everything (SGE): it rewards only successful agents and distributes the entire budget among them. In particular, contracts such as piece-rate or bonus-pool contracts, which under certain outcomes distribute only a fraction of the budget to successful agents, can never be optimal in this setting. Intuitively, the principal should be able to improve upon contracts that leave money on the table by using that leftover money to increase the rewards of some successful agent, thus incentivizing them to exert greater effort. [^1] This intuition is slightly misleading, since equilibrium forces imply that there isn’t necessarily an obvious way to distribute this left-over budget and induce higher effort from all agents. Nevertheless, we show that there is always some way to do this, and this implies that any optimal contract must be SGE. We further show that any SGE contract is optimal for some monotone objective of the principal. In other words, the maximal set of implementable equilibria is exactly the set of equilibria of SGE contracts.
@@
Our main contribution is then to introduce a subclass of SGE contracts, called Luce contracts, and to show that there is always a Luce contract that is optimal. A Luce contract is defined by assigning a weight to each agent and a (weak) priority order over the set of agents, so that the entire budget is distributed among the highest-priority successful agents in proportion to their weights. We show that Luce contracts are minimally sufficient to implement the maximal set, and thus, the principal can always optimize over Luce contracts. This result provides for a significant reduction in the complexity of the design problem, narrowing the search space from a $\Theta(n2^{n})$-dimensional space of complex and difficult-to-interpret contracts to an $(n-1)$-dimensional space of simple and easily interpretable contracts. We apply our results to the special case of two agents with quadratic costs, where our analysis uncovers an interesting robustness in the structure of the optimal contract to the heterogeneity in the agents’ cost parameters.
@@
There is a vast literature on principal-agent problems under moral hazard. In the canonical model with a single agent (Holmström (1979); Grossman and Hart (1992); Mirrlees (1976)), the principal offers a wage contract that defines the agent’s payment as a function of its observed output, and then the agent chooses some unobserved action (effort) that determines the distribution over outputs. A key finding is that optimal contracts reward the agent for output realizations that are informative about the target level of effort (informativeness principle), and may therefore be non-monotone in output. There has since been significant work studying variants of this single-agent model incorporating flexible actions (Georgiadis, Ravid, and Szentes (2024)), multiple tasks (Holmstrom and Milgrom (1991); Bond and Gomes (2009)), bounded payments (Jewitt, Kadan, and Swinkels (2008)), combinatorial actions (Dütting, Ezra, Feldman, and Kesselheim (2022)), and informationally robust design (Carroll (2015); Zhang (2023)). See Georgiadis (2022); Holmström (2017) and Lazear (2018) for recent surveys of this literature.
Our paper contributes to the literature studying a principal contracting with multiple agents. In these settings, agents may be rewarded not only based on their absolute performance but also on their relative performance or even the joint performance of the group. When agents’ observed outputs are subject to a common shock or involve subjective evaluations, the optimal contract may involve pay for relative performance (Green and Stokey (1983); Lazear and Rosen (1981); Malcomson (1986, 1984); Mookherjee (1984); Nalebuff and Stiglitz (1983); Imhof and Kräkel (2014)). And joint performance evaluation may be optimal in the presence of negatively correlated outputs (Fleckinger (2012)), complementarities in production (Alchian and Demsetz (1972)), incentives to help (Itoh (1991)), and unknown production environment (Kambhampati (2024)). Other related work on multi-agent contract design explores moral hazard in teams (Holmstrom (1982); Winter (2004); Battaglini (2006); Babaioff, Feldman, Nisan, and Winter (2012); Halac, Lipnowski, and Rappoport (2021); Dai and Toikka (2022)), computational challenges in finding optimal contracts (Castiglioni, Marchesi, and Gatti (2023); Dütting, Ezra, Feldman, and Kesselheim (2023); Ezra, Feldman, and Schlesinger (2024)), and design problems tailored to more specific environments (Halac, Kartik, and Liu (2017); Haggiag, Oren, and Segev (2022); Baiman and Rajan (1995)). In contrast to this literature, our work focuses on a principal contracting with multiple independent agents, where standard single-agent contracts would be optimal but are rendered infeasible due to budgetary constraints on the principal. [^2] There is also a related literature in contests, where the contract specifies a distribution of the budget based on how agents rank in terms of their output (Fang, Noe, and Strack (2020)). However, in our model, the principal observes only whether agents succeed or fail, not their exact effort levels or rankings. Thus, the reward schemes typically analyzed in contest settings are not applicable to our framework.
@@ 2 Model
There is a principal and $n$ risk-neutral agents. Each agent $i\in[n]$ is assigned an independent task, in which it may succeed or fail. The agents choose how much effort to exert towards succeeding in their respective tasks, captured by their choice of success probabilities: if agent $i$ chooses success probability $p_{i}\in[0,1]$, it incurs a cost $c_{i}(p_{i})$ in doing so, where $c_{i}:[0,1]\to\mathbb{R}$ is differentiable, strictly increasing, and strictly convex, with $c_{i}^{\prime}(0)=0$. The cost functions are common knowledge. [^3] We note that our model can also be interpreted in a more general framework but focusing on a restricted class of target-based contracts. Specifically, suppose each agent $i\in[n]$ may choose an effort level $e_{i}\in\mathbb{R}_{+}$, which induces a probability distribution over a set of observable outcomes $X_{i}$. Consider contracts in which the principal specifies a target $t_{i}\in X_{i}$ for each agent so that the agent is deemed to have succeeded if the realized outcome $\geq t_{i}$ and failed otherwise. Under standard assumptions, such as $X_{i}$ being ordered and higher effort inducing distributions that first-order stochastically dominate those under lower effort, our model can be interpreted as analyzing such target-based contracts within this general setting.
The principal would like to incentivize the agents to maximize the probability that they succeed in their tasks. Formally, we assume that the principal’s preference over success probability profiles is represented by a continuous, strictly increasing objective function $V(p_{1},\dots,p_{n})$. For example, if the principal obtains a profit of $w_{i}$ if agent $i$ succeeds in its task, and it cares about maximizing expected profit, its preference may be represented by the objective function
@@
Alternatively, if the principal gets a constant positive payoff as long as there is at least one successful agent, and 00 otherwise (as may be the case in a crowdsourcing for innovation example), its preference may be represented by
$V(p)=1-\Pi_{i=1}^{n}(1-p_{i}).$
In order to incentivize the agents, the principal can design a contract mapping observed outcomes to a reward for each agent. We assume that the principal has an exclusive-use budget $B>0$ that can be used for the purpose of rewarding the agents, and the principal gets no value from unused portions of the budget. Formally, the principal commits to a contract which specifies the fraction of the budget each agent receives depending on the observed outcome.
@@
$u_{i}(p)=\mathbb{E}_{p}[f_{i}(S)\cdot B]-c_{i}(p_{i})=\sum_{S\subset[n]}{\Pr}^{[n]}_{p}(S)\cdot f_{i}(S)\cdot B-c_{i}(p_{i})$
@@
Notice that the expected reward in linear in $p_{i}$, and since the cost is convex, the utility is strictly concave in $p_{i}$. It thus follows that pure-strategy Nash equilibrium exists. We will denote by $E:\mathcal{F}\rightrightarrows[0,1]^{n}$ the equilibrium correspondence, so $E(f)$ is the set of pure-strategy equilibria for the contract $f$ and $E^{-1}(p)$ is the set of contracts for which $p$ is an equilibrium. We further denote by $\mathcal{E}$ the set of implementable profiles:
@@
The principal’s problem is to choose a contract $f$ that induces an equilibrium $p\in E(f)$ that maximizes its objective $V$. As is standard, we assume that when a contract has multiple equilibria, the principal can select its most preferred equilibrium. [^4] For an illustration of multiplicity, suppose $B\geq\sum_{i\in[n]}c_{i}^{\prime}(1)$ and consider the bonus-pool contract under which agent $i$’s reward is $c_{i}^{\prime}(1)$ if all agents succeed and 00 otherwise. For this contract, it is easy to verify that both $p=(1,1,\dots,1)$ and $p=(0,0,\dots,0)$ are Nash equilibria. That said, if the budget is sufficiently small and the cost functions are sufficiently convex, we can use the diagonally strict concavity condition of Rosen (1965) to show that every contract has a unique equilibrium. Hence, the principal’s problem is to maximize its objective $V$ over the set of implementable profiles $\mathcal{E}$. As we show, $\mathcal{E}$ is compact, and since $V$ is strictly increasing, any maximizer of $V$ must be a maximal element of $\mathcal{E}$ with respect to the standard coordinatewise ordering. We denote by $\mathcal{P}$ the set of maximal equilibrium profiles in $\mathcal{E}$. Formally,
@@
Notice that if the budget is sufficiently large ($B\geq\sum_{i\in[n]}c_{i}^{\prime}(1)$), the principal can induce any effort profile $q\in[0,1]^{n}$ using the following piece-rate contract:
@@
Indeed, agent $i$’s expected utility in this case is $p_{i}\cdot c_{i}^{\prime}(q_{i})-c_{i}(p_{i})$ which is uniquely maximized at $p_{i}=q_{i}$. It follows that $\mathcal{E}=[0,1]^{n}$, $\mathcal{P}=\{(1,1,\dots,1)\}$, and the problem is trivial.
@@
Assumption 1 ensures that in equilibrium, no agent chooses $p_{i}=1$. We note that our results extend to situations with a larger budget as long as we restrict attention to equilibria that are interior.
3 Main results
In this section, we identify and characterize a family of contracts whose equilibria coincide precisely with the set of maximal equilibria.
Consider any contract $f\in\mathcal{F}$. At any effort profile $p\in[0,1]^{n}$, agent $i$’s utility can be conveniently expressed as
$u_{i}(p)=\big{(}\mathbb{E}_{p}[f_{i}(S)\,|\,i\in S]\cdot p_{i}+\mathbb{E}_{p}[f_{i}(S)\,|\,i\notin S]\cdot(1-p_{i})\big{)}\cdot B-c_{i}(p_{i}).$ (1)
Since $u_{i}$ is strictly concave in $p_{i}$, agent $i$ has a unique best response characterized by the first-order condition:
$c_{i}^{\prime}(p_{i})=\max\{0,\big{(}\mathbb{E}_{p}[f_{i}(S)\,|\,i\in S]-\mathbb{E}_{p}[f_{i}(S)\,|\,i\notin S]\big{)}\cdot B\}.$ (2)
We denote this unique best response by $b_{i}(f,p_{-i})$.
Equation (2) makes clear that rewarding agents when they fail reduces their incentive to exert effort, and suggests that the principal should avoid using contracts that reward unsuccessful agents. We introduce this class of contracts formally: